Probability that a randomly choose student in a statistics class submits his/her homework on time is 0.14. Each student does homework independently. In a statistics class of 50 students, what is the probability that at least 5 will do their homework on time?
Answer #1

solusian Here we have to find P(x 35) using Binomial P(225) > 1-placc5) Pfx 25) = 1-[plu=4) + P(x=3) +p fx = 2) +Pfx =)+P(x=0